# EMERGE: Integrating RAG for Improved Multimodal EHR Predictive Modeling

### v1

@@ 6.2.1 Comparing Different Modality Fusion Strategies
Table 4: Ablation study results of 1) comparing each modality with RAG enhancement, and 2) comparing different multimodal fusion networks. Bold and Underlined indicates the best and 2nd best performance. All metrics are multiplied by 100 for readability purposes.
@@
6.2.2 Comparing Different Time-series Encoders
@@
Figure 7: AUROC performance of four time-series encoders in in-hospital mortality prediction and 30-day readmission prediction tasks on two datasets.
6.2.3 Comparing Different Text Fusion Approaches
@@
Figure 8: AUROC performance of different text fusion methods in in-hospital mortality prediction and 30-day readmission prediction tasks on two datasets.
6.2.4 Comparing Internal Design of Fusion Module
@@
Figure 9: AUROC performance of different internal designs of our proposed fusion module in in-hospital mortality prediction and 30-day readmission prediction tasks on two datasets.
6.3 Further Analysis
6.3.1 Sensitivity to Hidden Dimensions and Learning Rates
To assess the sensitivity of our EMERGE framework to different hidden dimensions and learning rates, we conducted experiments on the MIMIC-III and MIMIC-IV datasets, as shown in Figure 10. The results indicate that a hidden dimension of 128 and a learning rate of 1e-3 yield the best performance. The minimal variations across different settings demonstrate EMERGE’s low sensitivity to these hyperparameters.
@@ 6.3.1 Sensitivity to Hidden Dimensions and Learning Rates
Figure 10: AUROC performance of sensitivity analysis to hidden dimensions (left) and learning rates (right) in in-hospital mortality prediction and 30-day readmission prediction tasks on MIMIC-IV.
6.3.2 Robustness to Data Sparsity
@@
Figure 11: AUROC performance across 5 Training Set Percentage in in-hospital mortality prediction (left) and 30-day readmission prediction (right) task on MIMIC-IV.
6.3.3 Case Study of Generated Summary
We include a case study in Figure 12, showcasing a comprehensive summary of the patient’s health status with EMERGE. It highlights a “high fraction of inspired oxygen”, which aligns with “fraction inspired oxygen too high” derived from time-series data, thus enabling further risk prediction. Disease entities such as adenocarcinoma, non-small cell lung cancer (NSCLC stage IV), among others, are all noted in the patient’s clinical records, demonstrating that the LLM meticulously follows the original data during the generation process. Additionally, the model conducts an analysis of the aforementioned diseases and concludes that there is a “complex medical history and the presence of multiple comorbidities”. It also forecasts a “high risk for in-hospital mortality, potential 30-day readmission” along with the auxiliary recommendation of “close monitoring and management”.
Figure 12: Prompt template to generate the patient health status summary (above) and the corresponding generated summary output (below).
7 Conclusions

### v2

@@ 6.2.1. Comparing Different Modality Fusion Strategies
Table 3. Ablation study results of 1) comparing each modality with RAG enhancement, and 2) comparing different multimodal fusion networks. Bold and Underlined indicates the best and 2nd best performance. All metrics are multiplied by 100.
@@
6.2.2. Comparing Different Time-series Encoders
@@
Figure 7. AUROC performance of four time-series encoders in in-hospital mortality prediction and 30-day readmission prediction tasks on two datasets.
6.2.3. Comparing Different Text Fusion Approaches
@@
Figure 8. AUROC performance of different text fusion methods in in-hospital mortality prediction and 30-day readmission prediction tasks on two datasets.
6.2.4. Comparing Internal Design of Fusion Module
@@
Figure 9. AUROC performance of different internal designs of our proposed fusion module in in-hospital mortality prediction and 30-day readmission prediction tasks on two datasets.
6.3. Further Analysis
6.3.1. Sensitivity to Hidden Dimensions and Learning Rates
To assess the sensitivity of our EMERGE framework to different hidden dimensions and learning rates, we conducted experiments on the MIMIC-III and MIMIC-IV datasets (Figure 10). The results indicate that a hidden dimension of 128 and a learning rate of 1e-3 yield the best performance. The minimal variations across different settings demonstrate EMERGE’s low sensitivity to these hyperparameters.
@@ 6.3.1. Sensitivity to Hidden Dimensions and Learning Rates
Figure 10. AUROC performance to various hidden dimensions (left) and learning rates (right) in in-hospital mortality and 30-day readmission prediction tasks on MIMIC-IV.
6.3.2. Robustness to Data Sparsity
@@
Figure 11. AUROC performance across 5 Training Set Percentage in in-hospital mortality prediction (left) and 30-day readmission prediction (right) task on MIMIC-IV.
7. Conclusions
@@
Acknowledgements.
This work was supported by the Defense Industrial Technology Development Program under Grant JCKY2021601B104.